LASER SPECTROSCOPIC CHARACTERIZATION OF THE GROUND X̃ ^{1}Σ_{g}^{+} AND EXCITED Ã ^{1}B_{1} STATES OF JET-COOLED LIOLI (”LITHIATED WATER”)

\underline{D. BELLERT}, DARIN W. WINN, and W.H. BRECKENRIDGE, Department of Chemistry, University of Utah, Salt Lake City, Utah 84112.

Various laser based spectroscopies, including R2PI (Resonance Two-Photon Ionization), LIF (Laser-induced Fluorescence), DF (Dispersed Fluorescence), and SEP (Stimulated Emission Pumping) have been utilized to determine several vibrational and rotational constants for the ground (X̃ ^{1}Σ_{g}^{+}) and first singlet excited state (Ã ^{1}B_{1}) of ^{7}Li^{16}O^{7}Li. From analysis of SEP spectra to (ν_{1}, ν_{2}^{ℓ}, ν_{3} = 0) vibrational states, the ^{7}Li^{16}O^{7}Li ground state is confirmed to be linear, with a bond length (R_{0}) of 1.611 \underline{+} 0.005 Å. The ^{1}B_{1} excited state is bent, with a bond angle of approximately 105^{∘}, and a bond length (R_{0}) of approximately 1.84 Å. These geometries are consistent with ab initio estimates, and with the idea that the bonding in both these states is essentially ionic (Li^{+1}O^{−2}Li^{+1} and Li^{+1/2}O^{−1}Li^{+1/2}, respectively).